Pip update
This presentation is the property of its rightful owner.
Sponsored Links
1 / 16

PIP Update PowerPoint PPT Presentation


  • 141 Views
  • Uploaded on
  • Presentation posted in: General

PIP Update. May 23 2012. Agenda. Summary Update Activities Budget Plans Update Slides Solid StateTom ModulatorsRene High Power Ferrite Testing Dennis. Update (PIP – Shutdown). RFQ Energy Problem Continues (multiple tests – studies) Schempp is coming to FERMI after IPAC

Download Presentation

PIP Update

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Pip update

PIP Update

May 23 2012


Agenda

Agenda

  • Summary Update

    • Activities

    • Budget

    • Plans

  • Update Slides

    • Solid StateTom

    • ModulatorsRene

    • High Power Ferrite Testing Dennis


Update pip shutdown

Update (PIP – Shutdown)

  • RFQ

    • Energy Problem Continues (multiple tests – studies)

    • Schempp is coming to FERMI after IPAC

    • Buncher pumping down

  • Cavity 7 Testing

    • Two tuners are off

    • Testing may start next week

  • LCW – Set to go

  • Booster RF – lots going on (slides later)

  • Bias Supply work – ready to purchase transformers

  • Linac High Power Write Up – on going

  • Linac Modulator – met with Continental


Budget budget update

Budget – Budget Update


Budget update

Budget Update


Budget continued

Budget Continued


Budget continued1

Budget Continued


Solid state driver upgrade update

Solid State Driver Upgrade Update

  • 1kW Solid State Amplifiers

    • 67 have been tested and calibrated.

    • 2 are in the queue to be tested. Completely assembled.

    • 1 was being tested, but needs a heat sink repair.

      • Temperature of channel A was 20°F higher than the rest.

    • 3 have combiners in them, need to finish assembly.

    • 8 chassis built, need heat sinks.

    • 3 chassis need to be built, add heat sinks to those.


Solid state driver upgrade update1

Solid State Driver Upgrade Update

  • Issues

    • Not too many issues with testing and calibrating.

      • Most common problem is usually an inability to match the currents to within 100mA per channel at 16dBm RF input and the output voltage at 0dBm RF input within ±1V.

        • Alleviated by swapping out Driver/Output FET’s.

    • 2 have had heat sink problems.

      • Improper cooling on one channel.

    • Miscellaneous minor issues (bad IC or capacitor, etc.)

    • No real labor issues for now. A little behind schedule but amp testing and calibration won’t be a showstopper.


Booster modulator upgrade

Booster Modulator Upgrade

  • All parts have been ordered

  • We have 6 part time technicians making subassemblies and 3 full time technicians assembling the Modulators.

  • There are 8 Modulators of the first batch of ten which are 25 % assembled. The other 12 Modulator are about 10 % assembled.

  • The difficult part is the wiring. Expect the wiring to be done by the end of August.

René Padilla, Matt Domeier


Continued

continued

  • After the wiring installing the equipment and hooking them up should be quit easy. Maybe 2 months work.

René Padilla, Matt Domeier


High power testing of ferrite cores for new booster ferrite bias tuners proton improvement plan

High power testing of ferrite cores for new Booster ferrite bias tuners(Proton Improvement Plan)

Dennis Barak

On behalf of John Reid, Matt Slabaugh, and Steve Hays


What has been completed

What has been completed:

  • Ferrite core samples have been obtained

  • Low power testing of samples has been completed by Ralph Pasquinelli (Nov. 2011)

    • See THIS document for more details

  • High power testing needs to be completed

  • The high power testing fixture has been refurbished mechanically

  • A high current PS has been designed and built by Steve Hays and co.

    • Capable of providing up to 30-kA


Pip update

Ferrite testing fixture


Pip update

High Current PS


We are currently working on

We are currently working on:

  • Finish preparing the high current PS at the testing location (E48 kicker room)

  • Connecting the high current PS to the testing fixture

  • Initial measurements and adjustments need to be made to the fixture

    • Measure resonant frequency of the structure

    • Correct the resonant frequency through capacitive loading

    • Measure unloaded and loaded Q-factor of structure

  • Complete high power test on each sample core


  • Login